OpenStudy (anonymous):

12x^2 - 4x -5 Factorize this completely plz. or any shortcut of breaking so, tell me.

OpenStudy (anonymous):

\[x=(-b \pm \sqrt{b^2 -4ac})/2a\] use this formula!

OpenStudy (gorv):

12x^2-10x+6x-5=0 2x(6x-5)+1*(6x-5)=0 (2x+1)(6x-5)=0 x=-1/2 or x=5/6

OpenStudy (anonymous):

where a is 12, b is -4, & c is -5. in the end u'll get to values for x, use them to get the final factors ie : if the two roots are r & t then : (x+r)*(x+t)

OpenStudy (anonymous):

 behold